OpenStudy (anonymous):

Which set of integers is not a Pythagorean triple? A. 9, 40, 41 B. 11, 60, 61 C. 15, 18, 21 D. 20, 21, 29

OpenStudy (mathstudent55):

Are you familiar with the Pythagorean Theorem?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

The should have a ratio of 3:4:5

OpenStudy (mathstudent55):

This is the part of it that is needed here. Here is a right triangle. The two sides that form the right angle have lengths a and b. They are called the legs. The side opposite the right angle has length c. It is called the hypotenuse. 

OpenStudy (mathstudent55):

In a right triangle, this equation is always true: \(a^2 + b^2 = c^2\) where a and b are the lengths of the legs, and c is the length of the hypotenuse.

OpenStudy (mathstudent55):

That equation is what we get from the Pythagorean Theorem.

OpenStudy (anonymous):

So how do I use that with this equation?

OpenStudy (mathstudent55):

The hypotenuse is always the longest side in a right triangle. That means the hypotenuse is always represented by c.

OpenStudy (mathstudent55):

For each group of three numbers, replace c with the largest number, and a and b with the other two, in either order. If the equation is true, it is a Pythagorean triple, and the triangle with those side lengths is a right triangle.

OpenStudy (mathstudent55):

Let's do the first choice as an example together. The numbers are 9, 40, 41 41 is the largest number, so we put in for c. 9 and 40 are used for a and b in either order. \(a^2 + b^2 = c^2 \) First, we just substitute the numbers in for a, b, and c. \(9^2 + 40^2 = 41^2\)

OpenStudy (mathstudent55):

Now we see if these three numbers make the equation true. 81 + 1600 = 1681 1681 = 1681 Since the equation is true (because 1681 really is equal to 1681), this is a Pythagorean triple. Now you can move on to the next choice.